OpenStudy (anonymous):

Tina is training for a biathlon. To train for the running portion of the race, she runs 10 miles each day, over the same course. The first 4 miles of the course is on level ground, while the last 6 miles is downhill. She runs 4 miles per hour slower on level ground than she runs downhill. If the complete course takes 1 hour, how fast does she run on the downhill part of the course?

OpenStudy (anonymous):

@Miracrown

OpenStudy (kropot72):

Let the downhill speed = s Then the level speed = s - 4 \[Time=\frac{distance}{speed}\] \[\frac{4}{s-4}+\frac{6}{s}=1 ............(1)\] Do you understand how I got equation (1)?

OpenStudy (anonymous):

yes, i do

OpenStudy (anonymous):

I multiply the left side by the right denominator & do the opposite the the right denominator, right?

OpenStudy (kropot72):

Good. Next we need to put the left hand side of (1) into a single fraction: \[\frac{4s+6(s-4)}{s(s-4)}=1\]

OpenStudy (anonymous):

\[\frac{ 4s+6s-24 }{ s^2 }\] right?

OpenStudy (anonymous):

s^2-4s, i mean

OpenStudy (anonymous):

Do I just add 4s+6s?

OpenStudy (kropot72):

Now by cross multiplying we get: \[4s+6s-24=s ^{2}-4s...........(2)\] Now rearrange (2) to get a quadratic to gthen solve for s.

OpenStudy (kropot72):

to then solve*

OpenStudy (anonymous):

I know the s^2 is first & it'll end with =0?

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

s^2-4s+6s-24=0, or do I have to change some signs?

OpenStudy (kropot72):

\[s ^{2}-14s+24=0............(3)\] Do you see where a sign change was needed?

OpenStudy (anonymous):

yes. now I factor it right?

OpenStudy (anonymous):

it's 12 & 2

OpenStudy (anonymous):

(both negatives

OpenStudy (anonymous):

(s-12)(s-4)

OpenStudy (anonymous):

i mean 2

OpenStudy (kropot72):

That is correct. Test each solution and discard the invalid one. Both possible solutions are positive,

OpenStudy (anonymous):

12 & 2 ! Thank you, can you help with the last one like this? please?

OpenStudy (kropot72):

If you plug the value s = 2 into equation (1) you get: \[\frac{4}{2-4}+\frac{6}{2}=-2+3=1\] This gives a negative value of time for the level part of the course. Therefore s = 2 is not a valid solution. Leaving 12 miles per hour as the solution for the downhill speed.

OpenStudy (anonymous):

That's what I put on my work.